b94h:

Pablo purchased adult and youth tickets for the football game. He bought x adult tickets for $74 each and y youth tickets for $35. Write an expression that can be used to show the total cost, C, of all tickets.

b94h:

Just started to use this app and i have done a quiz and got a perfect score!

b94h:



mxddi3:

Hi, welcome to QuestionCove!! seeing as our cost is C, we are going to have C= __x+___y so if x represents the number of adult tickets, and they each cost $74, what is going to be next to our x variable?

b94h:

it would be........

b94h:

C=74x+35y?

mxddi3:

yes, it would!! amazing job! do you understand how you got that answer?

b94h:

yes

mxddi3:

great! let me know if there's more i can help with. be sure to close the post when you are finished!

b94h:

Ella purchased a game that was on sale for 12% off. The sales tax in her county is 6%. Let y represent the original price of the game. Write an expression that can be used to determine the final cost of the game. And ok i will close a post when i am finished

mxddi3:

ok so we need to have y=____, right?

b94h:

yes

b94h:

yes we do need a y

mxddi3:

Let y represent the original price of a game. Ella purchased a game that was on sale 12% off which means that the new price of the game is 88% of the original price. follow me?

b94h:

yes

mxddi3:

\[y*0.88/100= 0.88y\] Next, the sales tax in her county is 6%, which means that the the now price is 106% to the original price, so: \[0.88y* 106= 0.88(1.06)y\] when you solve it out, your y is going to be the 88% and the 6% so you have .88y and 0.06y so y =− 0.88y +0.06y you want to subtract the 12 percent from the original and that's the 88 and then you add the 6 percent sales tax to it so the answer is y =− 0.88y +0.06y does this make any bit of sense or should i explain it differently?

b94h:

No no i makes perfect sense to me.

b94h:

it*

mxddi3:

ok, lol, if you want to double check it you can ask google or brainly

b94h:

Ok thank you for helping me

b94h:

You are a good helper

b94h:

:)

mxddi3:

tysm !

b94h:

Yw!

mxddi3:

when done, close the post! have a great day!

b94h:

Ok i will you do the same!